Circular No. 6111

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

PERIODIC COMET KOPFF (1994s) This comet has been recovered by C. W. Hergenrother, Lunar and

Planetary Laboratory, on CCD frames obtained by S. M. Larson with the 1.5-m reflector at the Catalina Station. The images are very faint (near threshold on the second night). The comet is stellar in appearance, with no hint of a coma or tail. The indicated correction to the prediction on MPC 22032 is ΔT = +0.02 day.

1994 UT  α_{2000}  δ_{2000}  m_{2}
Nov. 30.50175  10^{h}26^{m}18.^{s}01  +11^{o}35′17′′.7  22.8
30.50913  10 26 18.10  +11 35 17.7  22.0
30.51556  10 26 18.21  +11 35 16.5  23.0
30.52096  10 26 18.38  +11 35 16.5  21.8
Dec. 1.51258  10 26 36.35  +11 34 25.1  22.8
1.51769  10 26 36.56  +11 34 24.6  22.7
1.52237  10 26 36.64  +11 34 24.6  23.0

SUPERNOVA 1994ae IN NGC 3370 F. Patat, European Southern Observatory (ESO) and Padova Uni-

versity, reports that preliminary reduction of his CCD observations of SN 1994ae with the ESO/Dutch 0.9-m telescope at La Silla yields the following magnitudes: Dec. 1.36 UT, V = 13.12, V − R = −0.01; 3.35, 13.13, −0.01; 4.35, 13.13, −0.02. The constant luminosity over three days and the V − R colors suggest that SN 1994ae is at maximum light.

Visual magnitude estimates: Dec. 1.00 UT, 13.5 (Z. Vicián, Héhalom, Hungary); 1.08, 13.6 (L. Szentaskó, Veresegyház, Hungary).

SUPERNOVA 1994af IN ANONYMOUS GALAXY Patat also reports that a 10-min CCD exposure taken on Dec. 1.24 UT

with the ESO/Dutch 0.92-m telescope shows this object (cf. IAUC 6110) at magnitude R ≃ 18.04; on Dec. 2.16, the supernova appeared ∼ 0.1 mag fainter. SN 1994af was measured to be ∼ 3′′ east and 1′′ south of the galaxy’s nucleus; there are nearby stars of mag R = 18.36 and 17.05 that are located 100′′ east, 30′′ north, and 100′′ east, 90′′ south, of SN 1994af, respectively; a bright star appears 72′′ east and 98′′ north of the supernova.

1994 December 5 Daniel W. E. Green